OpenStudy (anonymous):

What is the slope Intercept form of the equation of the line in function notation? y+1/2=-2(x+3)

OpenStudy (mertsj):

Solve the equation for y

OpenStudy (anonymous):

y = -2x-5.5

OpenStudy (anonymous):

Pandae- thats not correct :(

jimthompson5910 (jim_thompson5910):

y+1/2=-2(x+3) y+1/2=-2x-6 y=-2x-6-1/2 y=-2x-13/2 So the equation of the line in slope intercept form is \[\Large y = -2x-\frac{13}{2}\]

OpenStudy (anonymous):

Correct in my world : D

jimthompson5910 (jim_thompson5910):

It should be y = -2x - 6.5 if you're representing it in decimal form